OpenStudy (anonymous):

The atomic weight of an element is 'a'.If this element occurs in nature as a triatomic gas,then the correct formula for the number of moles of gas in its 'w' g is:- a)\[\large \bf \frac{3w}{a}\] b)\[\large \bf \frac{w}{3a}\] c)\[\large \bf 3wa\] d)\[\large \bf \frac{a}{3w}\]

OpenStudy (anonymous):

@hartnn @ganeshie8

OpenStudy (anonymous):

Do we know what w stands for?

OpenStudy (anonymous):

yes, w stands for given mass/weight

OpenStudy (anonymous):

I thought that was a?

OpenStudy (anonymous):

no. a=atomic weight and w=given weight

OpenStudy (anonymous):

Are you adding multiple choice to it or something? You edited it and there is a) there now, so.

OpenStudy (anonymous):

here are the options(in the question)

OpenStudy (anonymous):

@halorazer

OpenStudy (anonymous):

I do believe that it is B. This is because on the periodic table, it gives the measurements for a single atom of something. So if we are trying to determine the amount of moles, for, say, ozone (O3.) And we had 48 grams of it, we would do: \[\frac{ 48 }{ 16 \times 3 } = \frac{ 48 }{ 48 } = 1 mole O3\]

OpenStudy (anonymous):

ohkk!!! thnx...

OpenStudy (anonymous):

48 being the given weight, 3 being because it's a triatomic molecule, and the 16 coming from the atomic weight of the oxygen atom.

OpenStudy (anonymous):

You are very welcome. :)